# STEM-PoM: Evaluating Language Models Math-Symbol Reasoning in Document Parsing

### v1

@@ Abstract
Advances in large language models (LLMs) have spurred research into enhancing their reasoning capabilities, particularly in math-rich STEM documents. While LLMs can generate equations or solve math-related queries, their ability to fully understand and interpret abstract mathematical symbols in long, math-rich documents remains limited. In this paper, we introduce STEM-PoM, a comprehensive benchmark dataset designed to evaluate LLMs’ reasoning abilities on math symbols within contextual scientific text. The dataset, sourced from real-world ArXiv documents, contains over 2K math symbols classified as main attributes of variables, constants, operators, and unit descriptors, with additional sub-attributes including scalar/vector/matrix for variables and local/global/discipline-specific labels for both constants and operators. Our extensive experiments show that state-of-the-art LLMs achieve an average of 20-60% accuracy under in-context learning and 50-60% accuracy with fine-tuning, revealing a significant gap in their mathematical reasoning capabilities. STEM-PoM fuels future research of developing advanced Math-AI models that can robustly handle math symbols.
@@ 1 Introduction
Large language models (LLMs) have demonstrated exceptional reasoning abilities across numerous fields [17, 13, 12, 25, 15]. With the increasing shift towards applying LLMs to complex tasks [6, 23, 39], the need for supplementary data beyond the general pre-trained datasets has become increasingly important. Among these, mathematical reasoning tasks [10, 18] have recently drawn the attention of several researchers [19, 2, 45, 28] (see Backgrounds in Appendix B). In particular, Part-of-Math Tagging [43], the mathematical analog to part-of-speech tagging [36] where mathematical tokens are classified according to a given taxonomy of attributes, continues to gain interest but lacks the foundational datasets that can support advanced NLP tasks [43, 38, 37]. In addition, integrating mathematical language into NLP models remains a substantial challenge [3, 29], especially in the realm of document parsing [8, 24, 44]. Traditional semantic parsing methods like LateXML [31] or arXMLiv [22] often fall short when applied to math-rich documents, where precision and structured syntax are paramount [14, 32, 41]. These methods struggle to accurately perform pattern matching between abstract mathematical symbols and their corresponding XML tag notations. Similarly, recent advanced LLMs, such as ChatGPT [26], also face difficulties in understanding and reasoning with abstract mathematical symbols due to their contextual polymorphism [35] (as Figure 3 shown). For example, in the linear equation: $y=mx+p$, $y$ is categorized as a variable. Whereas in the cross-entropy loss function: $\mathcal{L}(x,y)=-\sum_{i=1}^{N}x_{i}\log(y_{i})$, the symbol $y$ represents the fixed target labels, which is considered a constant for a given dataset. Without the corresponding contextual information of a mathematical symbol, LLMs are unable to distinguish between different attributes of the symbol and cannot effectively process related mathematical reasoning tasks. Thus, tagging math symbols within domain-specific contexts is essential for language models.
@@
In this paper, we introduce a novel benchmark dataset, STEM-PoM, designed to evaluate the reasoning capabilities of language models on mathematical symbols across different domains. The STEM-PoM dataset consists of 2,109 instances extracted from a random sampling of over 10,000 arXiv manuscripts, which are math-rich documents spanning domains such as Mathematics, Physics, Chemistry, and more. We provide a mathematical symbol for each dataset instance, its order in the document, its main and sub-level attributes, and the corresponding text or expression from the original arXiv paper containing the symbol. Each mathematical symbol in the dataset is classified according to two levels of attributes [42]. The first-level attribute categorizes the symbol as variable, constant, operator, or unit descriptor. The second-level attribute further classifies the symbol into one of six types based on its first-level category: scalar, vector, matrix, local, global, or discipline-specific. Figure 1 illustrates the dataset’s category distribution. To further enrich the STEM-PoM dataset with additional arXiv manuscripts and other math-rich document resources, we also design the STEM-PoM Labeler, a feasible method for assisting dataset generation by automatically searching, extracting, and recording hand-labeled mathematical symbols and their corresponding context from long-text documents.
@@
We conduct thorough experiments on the STEM-PoM dataset to assess the mathematical reasoning abilities of seven open- and closed-source language models, including LSTM [11], Mixtral-8x7B [20], Llama2-13B [40], Llama3-80B [9], Claude-3.5-sonnet [4], GPT-3.5, and GPT-4o [1] with various prompting and fine-tuning strategies. The experimental results indicate that STEM-PoM distinguishes the performance levels across different LLMs, offering insights into the mathematical symbol reasoning abilities of these models. In addition, we investigate and analyze the influence of context length on the ability of language models to understand mathematical symbols.
2 STEM-PoM Dataset
2.1 Data Annotation
Source of Mathematical Symbols. The first crucial step in constructing the dataset is selecting high-quality mathematical symbols. For STEM-PoM, we primarily collect these symbols from two sources: 1. Public math-symbol datasets, where we directly utilize candidate mathematical symbols from the mathematical token definition extraction benchmark, MTDE [14]. 2. Raw ArXiv papers [7], where we apply the STEM-PoM Labeler to identify and extract mathematical symbols from the ArXiv dataset. We include a detailed description of each source dataset in Appendix A.2.
Dataset Construction. After obtaining the mathematical symbols, we categorize each symbol into different attributes and assign corresponding information to construct the STEM-PoM dataset. Specifically, we first extract the file name and symbol order for each mathematical symbol. Then, for each symbol, we extract the contexts in which the symbol appears, using several predefined lengths. Following this, we manually classify each symbol into four primary categories: Variable, Constant, Operator, and Unit Descriptor. For Variable, Constant, and Operator, we further categorize them into sub-level categories. The variable is classified as Vector, Scalar, or Matrix, while Constant and Operator are categorized as Local, Global, or Discipline-Specific. Table 2 outlines the overall dataset structure. We manually examine each entry of the dataset thoroughly to ensure its robustness and correctness. We provide a detailed explanation of the dataset structure in Appendix A.3 and the definitions of each level’s attributes in Appendix A.4. Additionally, the STEM-PoM Labeler is described in Appendix A.5.
2.2 Dataset Statistics
@@ 3 Experiments
3.1 Setup
Models. To thoroughly evaluate our dataset across models with varying parameter sizes, we utilize the following models: LSTM framework [11], Llama-2-13B [40], Mixtral-8x7B-v0.1 [20], and GPT-3.5-turbo-0125 [1].
Evaluation Metrics. We apply the Precision Accuracy as our metric for the mathematical symbol classification task, the metric can be formulated as: Precision Accuracy=Number of correct predictionsTotal number of $\textit{Precision Accuracy}=\frac{\textit{Number of correct predictions}}{\textit{Total number of samples}}$
Training & Inference Details. We evaluate several models under both pre-training and fine-tuning settings. Specifically, we train an LSTM model with varying layers and apply the LoRA method [16, 47], a PEFT technique, to GPT-3.5. We evaluate other models in the in-context learning setting. Appendix C provides the training and model parameter details.
@@ 3.1 Setup
Claude3.5-Sonnet One Sentence 63.7% 58.6% 62.5% 65.7% 67.8%
@@
Table 3: First-level classification accuracy with various context lengths.
@@
(Vanilla) Scalar Vector Matrix Local DS Global Local DS Global
LSTM 13.8% 15.1% 17.2% 19.2% 17.8% 22.2% 16.6% 11.3% 14.6%
Llama2-13B 27.3% 24.4% 21.8% 33.6% 31.5% 33.6% 32.4% 28.3% 32.7%
Mistral-8x7B 36.9% 35.8% 21.6% 34.8% 31.2% 37.8% 36.4% 34.8% 35.7%
Llama3-80B 38.2% 34.1% 26.7% 37.6% 35.2% 36.1% 39.1% 32.3% 40.2%
Claude3.5-Sonnet 53.2% 49.7% 55.8% 55.9% 53.1% 49.6% 56.3% 52.2% 55.9%
GPT-3.5 44.5% 45.8% 48.3% 48.5% 42.9% 44.3% 48.4% 43.5% 49.7%
GPT-4o 54.6% 51.3% 58.6% 58.4% 54.1% 56.2% 60.5% 57.3% 58.5%
@@
3.2 First-Level Classification Results.
Table 3 presents the accuracy results for different models across varying context lengths. The result shows that the small-parameter-size language model such as the LSTM struggles with lower accuracy, achieving between 18.7% and 22.6%. In contrast, larger models, such as Llama2-13B and Mistral-8x7B, show marked improvements as context length increases, with Mistral-8x7B reaching up to 53.6% on the full manuscript. In addition, Claude3.5-Sonnet achieves comparable performance with GPT-4o across all context lengths, with accuracy consistently above 63.7% and up to 66.7%. GPT-based models exhibit stronger performance overall, with GPT-3.5 achieving between 56.8% and 60.6%. GPT-4o further improves across all context lengths, outperforming other models with an overall accuracy of 68.5% with the full manuscript input. The performance gap between models remains consistent as context length increases. For instance, GPT-4o outperforms Llama3-80B by 16.0%, 14.4%, and 16.8% for context lengths of one sentence, ten sentences, and the full manuscript, respectively. This consistent performance gap suggests that larger models with more pre-trained knowledge, such as GPT-4o and Claude3.5-Sonnet, exhibit superior scalability with longer contexts. These models can more effectively leverage extended context lengths to distinguish between mathematical symbols and other nuanced elements in the input prompts. On the other hand, the overall performance gain from increasing context length is more pronounced in smaller models, such as Llama2-13B and Mistral-8x7B, which have less pre-trained knowledge. These models benefit more from extended context as they rely on additional information to compensate for their limited pre-training. Larger models like GPT-4o and Claude3.5-Sonnet, which come with extensive pre-trained knowledge, show relatively smaller performance gains as context length increases.

### v2

@@ Abstract
Advances in large language models (LLMs) have spurred research into enhancing their reasoning capabilities, particularly in math-rich STEM (Science, Technology, Engineering, and Mathematics) documents. While LLMs can generate equations or solve math-related queries, their ability to fully understand and interpret abstract mathematical symbols in long, math-rich documents remains limited. In this paper, we introduce STEM-PoM, a comprehensive benchmark dataset designed to evaluate LLMs’ reasoning abilities on math symbols within contextual scientific text. The dataset, sourced from real-world ArXiv documents, contains over 2K math symbols classified as main attributes of variables, constants, operators, and unit descriptors, with additional sub-attributes including scalar/vector/matrix for variables and local/global/discipline-specific labels for both constants and operators. Our extensive experiments demonstrate that state-of-the-art LLMs achieve an average accuracy of 20–60% under in-context learning and 50–60% with fine-tuning, highlighting a substantial gap in their ability to classify mathematical symbols. By improving LLMs’ mathematical symbol classification, STEM-PoM further enhances models’ downstream mathematical reasoning capabilities. The code and data are available at https://github.com/jiaruzouu/STEM-PoM.
STEM-PoM: Evaluating Language Models Math-Symbol
Reasoning in Document Parsing
Jiaru Zou1thanks: Corresponding authors, Qing Wang1, Pratyush Thakur1, Nickvash Kani [^1] footnotemark: 1 1 University of Illinois Urbana-Champaign {jiaruz2, kani}@illinois.edu
@@ 1 Introduction
Figure 1: Total frequency (appearances) of the top-15 mathematical symbols in 1,000 randomly sampled ArXiv math-rich documents. Each math symbol contemporary belongs to multiple main attribute categories depending on its related context and mathematical expression. This illustrates the contextual polymorphism of a single math symbol.
Large language models (LLMs) have demonstrated exceptional reasoning abilities across numerous fields (Huang and Chang, 2022; Hadi et al., 2023, 2024; Li et al., 2024; Zheng et al., 2024). With the increasing shift towards applying LLMs to complex tasks (Brown, 2020; Kojima et al., 2022; Sui et al., 2023; Zou et al., 2024), the need for supplementary data beyond the general pre-trained datasets has become increasingly important. Among these, mathematical reasoning tasks English (2013); Ilany et al. (2010) have recently drawn the attention of several researchers Imani et al. (2023); Ahn et al. (2024); Zhang et al. (2024); Lu et al. (2023). In particular, Part-of-Math Tagging Youssef (2017), the mathematical analog to part-of-speech tagging Schmid (1994), where mathematical tokens are classified according to a given taxonomy of attributes, continues to gain interest with the integration of LLMs math reasoning.
However, despite this growing interest, Part-of-Math Tagging currently still lacks the foundational datasets that are crucial for supporting advanced NLP tasks Youssef (2017); Shan and Youssef (2021, 2024). In addition, integrating mathematical language into NLP models remains a substantial challenge Alshamari and Youssef (2020); Meadows and Freitas (2022), especially in the realm of document parsing Dridan and Oepen (2013); Lam et al. (2008); Zhang et al. (2019). Traditional semantic parsing methods such as LaTeXML Miller (2011) often fall short when applied to math-rich documents, where precision and structured syntax are paramount (Hamel et al., 2022; Paster et al., 2023; Wang et al., 2023). These methods struggle to accurately perform pattern matching between abstract mathematical symbols and their corresponding XML tag notations.
@@
Similarly, recent advanced LLMs, such as ChatGPT (Liu et al., 2023), also face difficulties in understanding and reasoning with abstract mathematical symbols due to their contextual polymorphism Ditchfield (1994); Fiore and Hamana (2013); Murase et al. (2023), as demonstrated in Figure 1. As an intuitive example, in the linear equation: $y=mx+p$, $y$ is categorized as a variable. Whereas in the cross-entropy loss function: $\mathcal{L}(x,y)=-\sum_{i=1}^{N}y_{i}\log(x_{i})$, the symbol $y$ represents the fixed target labels, which is considered a constant for a given dataset. Without the corresponding contextual information of a mathematical symbol, LLMs are unable to distinguish between different attributes of the symbol and cannot effectively process related mathematical reasoning tasks. Thus, tagging math symbols within domain-specific contexts is essential for language models.
@@
In this paper, we introduce a novel benchmark dataset, STEM-PoM, designed to evaluate the reasoning capabilities of language models on mathematical symbols across different domains. The STEM-PoM dataset consists of 2,109 mathematical token instances extracted from a random sampling of 10,000 arXiv manuscripts, which are math-rich documents spanning domains such as Mathematics, Physics, Chemistry, and more. For each dataset instance, we provide a specific mathematical symbol with the corresponding symbol order in the document, main and sub-level attributes, and the related text information from the original ArXiv paper. Each mathematical symbol in the dataset is classified according to two levels of attributes Wikipedia (2023). The first-level attribute categorizes the symbol as variable, constant, operator, or unit descriptor. The second-level attribute further classifies the symbol into one of six types based on its first-level category: scalar, vector, matrix, local, global, or discipline-specific. Figure 2 illustrates the category distribution of the dataset. To further enrich the STEM-PoM dataset with additional arXiv manuscripts and other math-rich document resources, we also design the STEM-PoM Labeler, a feasible toolkit to assist dataset generation by automatically searching, extracting, and recording hand-labeled mathematical symbols with their corresponding context from STEM articles.
We conduct extensive experiments on the STEM-PoM dataset to assess the mathematical reasoning abilities of a large amount of open- and closed-source vanilla language models. Our experimental results indicate that existing language models generally struggle with understanding mathematical symbols, even when provided with relevant context. Extending the context length to a certain degree and fine-tuning on symbol-related material can enhance their ability to classify mathematical symbols. We further examine the relationship between the model’s mathematical symbol classification and mathematical reasoning capabilities by evaluating its performance on challenging downstream tasks, such as OlympiadBench He et al. (2024), after enhancing its classification accuracy through fine-tuning on STEM-PoM. Last, we provide detailed dataset analysis and case studies to investigate the influence factors such as context length on the ability of language models to understand mathematical symbols. In summary, our main contributions are:
  • We propose a novel task combining Part-of-Math Tagging with document parsing to assess LLMs’ mathematical symbol classification abilities.
  • We introduce STEM-PoM, a benchmark dataset comprised of over 2,000 annotated instances extracted from math-rich documents spanning fields including Mathematics, Physics, and Chemistry. Each math symbol in the dataset is meticulously labeled with multiple hierarchical attributes, reflecting its role and context within the original document.
  • We evaluate LLMs’ mathematical reasoning on STEM-PoM, revealing performance variations across different models and highlighting the role of math symbol understanding in enhancing LLMs’ mathematical reasoning abilities.
@@
2 Backgrounds
Part-of-Math (PoM) Tagging. The part-of-math tagging task draws inspiration from similar tagging tasks such as part-of-speech tagging Schmid (1994). In the PoM context, the goal is to label individual mathematical tokens or expressions in math formulas with their corresponding mathematical roles. Youssef (2017) collected mathematical content, such as formula representation and tagging for specific mathematical formula translations and verifications, including converting formulae into semantic LaTeX or testing with tools like CAS (Computer Algebra Systems). However, this focus on structured and narrow formula translations does not align with the broader, more diverse text-based tasks required to assess NLP models, due to the lack of scalability features in the collected math symbols. Shan and Youssef (2021, 2024) recently evaluated the potential of leveraging LLMs for automated annotation and Part-of-Math tagging of math symbols conducted on the Digital Library of Mathematical Functions (DLMF) Lozier (2003). Since the source of math symbols is only one manuscript, the mathematical tokens collected only have a single classification type and are self-consistent. In contrast, our dataset incorporates the inherent messiness of published literature across several STEM subjects, where these domain-specific math symbols can have multiple classifications or meanings depending on the discipline and related contextual information.
Math Symbol Annotation. In the context of mathematical symbol annotations, a common approach is to simply extract definitions for mathematical tokens, transforming part-of-math tagging into a name-entity-recognition (NER) task Asakura et al. (2021); Shan and Youssef (2024). While NER does evaluate an AI model’s ability to process text, it is fundamentally different from part-of-math tagging, which requires the model to understand and classify mathematical tokens based on a specific taxonomy. Unlike NER, part-of-math tagging tasks defined in our dataset demand deeper comprehension of the mathematical symbol.
Large Language Models. Pre-trained large language models (LLMs) have become a cornerstone in modern NLP Rosenfeld (2000); Zhao et al. (2023). Early approaches were based on N-gram models, but with the advent of distributed word embeddings Bengio et al. (2000); Mikolov (2013), neural language models gained prominence. The scalability and performance improvements introduced by these models and the availability of vast textual data have enabled the unsupervised pre-training of LLMs. These models, often referred to as foundation models Radford et al. (2019); Kojima et al. (2022), can then be fine-tuned on smaller, task-specific datasets to adapt them for various downstream applications. For STEM-PoM, we apply one traditional sequence-based NLP model, LSTM Graves and Graves (2012), and several most updated LLMs for our dataset evaluation.
@@ 2 Backgrounds
3 STEM-PoM Dataset
In this section, we introduce our constructed benchmark dataset, STEM-PoM. First, we describe the source data used for extracting mathematical symbols and text information in Section 3.1. Next, we outline the data annotation process in Section 3.2 and present the dataset statistics in Section 3.2. Finally, we provide details on our dataset labeling tool STEM-PoM labeler in Section 3.4.
3.1 Source Dataset
The first crucial step in constructing the dataset is selecting high-quality mathematical symbols. For STEM-PoM, we primarily collect these symbols from two sources: 1. Public math-symbol datasets, where we directly utilize candidate mathematical symbols from the mathematical token definition extraction benchmark, MTDE (Hamel et al., 2022). 2. Raw ArXiv papers Clement et al. (2019), where we identify and extract mathematical symbols with corresponding context including math equations, math symbol definitions, and related text sentences from the dataset. The detailed description of each raw dataset is provided below:
MTDE
Hamel et al. (2022) contains approximately 10,000 entries of mathematical symbol names along with their defined contexts. Each entry includes a ’short’ definition and a ’long’ definition. A short definition is a single-word definition, while a long definition consists of one or more words. The data was collected through random sampling from mathematical and scientific arXiv preprint manuscripts, covering a broad range of disciplines such as Physics, Computer Science, and Biology. Note that all math symbols in the MTDE dataset are pre-filtered and well-defined by the authors, ensuring our annotation process minimizes the risk of misinformation or repetition issues. In our dataset pre-processing, we also make sure the candidate data is generated via a corpus crawler and subsequently pruned and cleaned manually.
ArXiv Paper Dataset
Kohlhase et al. (2024) contains 1.7 million arXiv articles, spanning a wide range of disciplines, including Mathematics, Physics, Chemistry, Economics, and Computer Science. We will provide a detailed explanation of how STEM-PoM is constructed from the raw source dataset in the next subsection.
3.2 Dataset Construction
To construct our dataset, we randomly sampled 10,000 articles from the ArXiv Paper Dataset across various domains. We manually verified that each manuscript is math-rich, containing numerous mathematical expressions and symbols. Using pre-filtered mathematical symbols from the MTDE, we engaged human-expert annotators to utilize the STEM-PoM labeler to: (1) align each mathematical symbol with its source paper, (2) identify relevant contextual information for each symbol, and (3) address disambiguation and aliasing by annotating all possible meanings for symbols with multiple interpretations and specifying the context in which each meaning applies. Through the rigorous selection process, 453 of the 10,000 articles are identified as the final source articles with high-quality contextual matches for the symbols. Each article contains an average of 4.7 annotated mathematical symbols (2109 in total).
After obtaining the mathematical symbols, we categorize each symbol into different attributes and assign surrounding context information to construct the STEM-PoM dataset. Specifically, we first extract the file name and symbol order for each mathematical symbol. Then, for each symbol, we extract the contexts in which the symbol appears, using several predefined lengths. Following this, we manually classify each symbol into four main attribute (first-level) categories: Variable, Constant, Operator, and Unit Descriptor. For the Variable, Constant, and Operator, we further categorize them into sub-attribute (second-level) categories. A variable is classified as a Vector, Scalar, or Matrix, while a Constant or Operator is categorized as Local, Global, or Discipline-Specific. Table 2 outlines the overall dataset structure. We manually examine each entry of the dataset thoroughly to ensure its robustness and correctness. The overall dataset pipeline is demonstrated in Figure 2. We also provide a detailed explanation of the dataset structure in Appendix A.1 and the definitions of each level’s attributes in Appendix A.2.
Dataset Statistics
3.3 Dataset Quality Control
During the dataset construction, we manually evaluate the quality and applicability of the annotated data. Specifically, we process the evaluation process through both consistency checks and Inter-annotator agreements.
Consistency Checks.
To ensure reliability, we perform the following consistency checks: each article is assigned to domain-specific experts for labeling, and we provide precise definitions and examples for each label and category to guide the process. After labeling, we randomly select a subset of the labeled data for consistency evaluation. Additionally, we analyze data points with inconsistent labels, facilitate discussions among annotators to resolve ambiguities and reach a consensus, and refine the annotation guidelines to address any common sources of confusion.
Inter-annotator Agreements.
Regarding the IAA in our human annotation process, we engage 33 domain experts as our annotators to annotate domain-specific articles. After the initial labeling process, the rest annotators (5 Annotator/Discipline on average) from the same domain reviewed the annotated data by switching their labeling components. For the Inter-annotator Agreement, we measure with Cohen’s Kappa value and ensure the value ranges from 0.81 to 1 (0.903 on average across all labeled data), regarding the dataset’s mathematical symbols and their corresponding attributes. To maintain accuracy, annotators have access to the original papers through the STEM-PoM labeler and annotate the attributes of the data based on the content of the original papers. This thorough review process reinforces the reliability of our annotated dataset.
3.4 STEM-PoM Labeler
We developed a toolkit named STEM-PoM Labeler to assist human experts in retrieving math symbols along with their corresponding article contexts, annotating the dataset, and conducting inter-annotator consistency checks to ensure agreement in dataset labeling. For a detailed description of the toolkit’s design, please refer to Appendix A.3.
@@ 4 Experiments
Claude3.5-Sonnet Anthropic (2024) One Sentence 63.7% 58.6% 62.5% 65.7% 67.8%
@@
Table 3: First-level classification accuracy with various context lengths. Here. One sentence/Ten Sentences/Full Manuscript refers to the context size of completed sentences as contextual information for each math symbol.
@@
Scalar Vector Matrix Local DS Global Local DS Global
LSTM Graves (2013) 13.8% 15.1% 17.2% 19.2% 17.8% 22.2% 16.6% 11.3% 14.6%
Llama2-13B Touvron et al. (2023) 27.3% 24.4% 21.8% 33.6% 31.5% 33.6% 32.4% 28.3% 32.7%
Mistral-8x7B Jiang et al. (2024) 36.9% 35.8% 21.6% 34.8% 31.2% 37.8% 36.4% 34.8% 35.7%
Llama3.1-70B Dubey et al. (2024) 38.2% 34.1% 26.7% 37.6% 35.2% 36.1% 39.1% 32.3% 40.2%
Claude3.5-Sonnet Anthropic (2024) 53.2% 49.7% 55.8% 55.9% 53.1% 49.6% 56.3% 52.2% 55.9%
GPT-3.5-turbo Achiam et al. (2023) 44.5% 45.8% 48.3% 48.5% 42.9% 44.3% 48.4% 43.5% 49.7%
GPT-4o Hurst et al. (2024) 54.6% 51.3% 58.6% 58.4% 54.1% 56.2% 60.5% 57.3% 58.5%
@@
4.1 Setups
Models.
To thoroughly evaluate our dataset across models with varying parameter sizes, we utilize the following models: LSTM Graves (2013), Mixtral-8x7B-v0.1 Jiang et al. (2024), Llama2-13B Touvron et al. (2023), Llama3.1-70B Dubey et al. (2024), Claude-3.5-Sonnet Anthropic (2024), GPT-3.5-Turbo-0125111https://platform.openai.com/docs/models/gpt-3-5-turbo Achiam et al. (2023), and GPT-4o-2024-08-06222https://platform.openai.com/docs/models/gpt-4o Hurst et al. (2024).
Evaluation Metrics.
We apply the Precision Accuracy as our metric for the mathematical symbol classification task, the metric can be formulated as:
Precision Accuracy=Number of correct predictionsTotal number of $\textit{Precision Accuracy}=\frac{\textit{Number of correct predictions}}{\textit{Total number of samples}}$
Training & Inference Details.
We evaluate several models under both pre-training and fine-tuning settings. Specifically, we train an LSTM model with varying layers and apply the LoRA method, a parameter-efficient fine-tuning (PEFT) technique, to several updated LLMs. We also evaluate additional LLMs under the in-context learning setting. The training and model parameter details are provided in Appendix B.
Models (fine-tuned) Context Length Overall Variable Constant Operator Unit Descriptor
Llama-2-13B Touvron et al. (2023) Ten Sentences 45.5% 38.4% 41.7% 50.6% 51.3%
Mixtral-8x7B Jiang et al. (2024) Ten Sentences 55.7% 46.3% 51.9% 64.2% 60.3%
Llama3.1-70B Dubey et al. (2024) Ten Sentences 62.4% 56.6% 54.5% 70.8% 67.5%
GPT-3.5-turbo Achiam et al. (2023) Ten Sentences 66.9% 65.4% 66.6% 71.3% 64.5%
GPT-4o Hurst et al. (2024) Ten Sentences 70.3% 71.3% 74.1% 75.2% 68.1%
Table 5: First-level classification accuracy with various context lengths. Here. "One sentence/Ten Sentences/Full Manuscript" refers to the context size of completed sentences used as contextual information for testing each mathematical symbol.
4.2 First-Level Classification Results
Table 3 summarizes the accuracy results of various models across different context lengths. To extract context information, we implement a surrounding window provided for domain human expert annotators to meticulously select the most relevant completed sentences surrounding the math symbol. Through the careful context selection process, we ensure that the information included in the prompts is accurate and contextually relevant. The result shows that the small-parameter-size language model such as the LSTM struggles with lower accuracy, achieving between 18.7% and 22.6%. In contrast, larger models, such as Claude3.5-Sonnet and GPT-4o, show marked improvements as context length increases, with accuracy consistently above 63.7% and up to 73.8%. We also found that the performance gap between models remains consistent as context length increases. To demonstrate, GPT-4o outperforms Llama3.1-70B by 16.0%, 14.4%, and 16.8% for context lengths of one sentence, ten sentences, and the full manuscript, respectively. This consistent performance gap suggests that larger models with more pre-trained knowledge, such as GPT-4o, exhibit superior scalability with longer contexts.
Another notable observation is that the overall performance gain from increasing context length is more pronounced in smaller models, such as Llama2-13B and Mistral-8x7B, which have less pre-trained knowledge. These models benefit more from extended context as they rely on additional information to compensate for their limited pre-training. Larger models like GPT-4o and Claude3.5-Sonnet, which come with extensive pre-trained knowledge, show relatively smaller performance gains as context length increases.